# Variability in the massive black hole binary candidate SDSS J2320+0024: no evidence for periodic modulation

### v1

@@ 3.2 Fit and model selection
Since our goal is to estimate the Bayesian evidence of the proposed models, we adopted a nested sampling algorithm for the parameter inference (Skilling, 2004). In particular, following what we did in previous works (e.g., Rigamonti et al., 2022, 2023), we employed the python-based implementation raynest (Veitch et al., 2017), which efficiently computes the marginal-likelihood (Eq. 1) using the GPyTorch package (see Gardner et al., 2021) [^5] While Eq. (1) is a marginal-likelihood for a single GP, it is instead to be interpreted as a likelihood when exploring GPs with different parameters.. We performed a parameter estimation with 1000 live points, assuming log-uniform priors for all the parameters, except for $\gamma$, which is assumed to have a uniform prior. The prior ranges adopted in all the models are reported in Tab. 1.
@@
Figure 1: Best-fit for the g-band light-curve of J2320+0024. The black points represent the data with their errors, the blue solid line and the shaded area represent the median, 16th, and 84th percentiles from the posterior distributions, and the red dashed line refers to the maximum likelihood model. From top to bottom: DRW (a), modified DRW (b), QPO (c), and Periodic (d), we refer to Tab. 1 for the estimated best-fit parameters.
@@
Fig. 1 shows the best-fit light-curves with the associated errors for the four models discussed in Sec. 3.1: DRW (a), Modified DRW (b), QPO (c), and Periodic (d). In all the plots, the black points represent the data with their associated uncertainties; the solid blue curves and their associated blue shaded area refer to the median, 16th, and 84th model percentiles computed propagating the full shape of the posterior distribution [^6] For each set of kernel parameter combination drawn from the posterior distribution, we sample 100 realizations from the GP and we take the median, 16th, and 84th among all the realization on all the set of sampled kernel parameters.; and, finally, the red dashed curve refers to the model that maximizes the likelihood. The best-fit model parameters (either the percentiles or the likelihood maxima) together with their corresponding Bayesian evidence values are reported in Tab. 1. It is important to clarify that while the maximum likelihood models (i.e., red dashed lines) are obtained using the ”Maximum likelihood” parameters reported in Tab. 1, this is not the case for the median, 16th, and 84th models. Indeed, in agreement with our Bayesian approach, the blue lines and the shaded areas presented in Fig. 1 are the median, 16th, and 84th percentiles computed after evaluating the GP model on all the parameter combinations describing the posterior distribution. Fig. 2 displays a comparison of the marginalised posterior distribution of the DRW (blue), Modified DRW (red), QPO (green), and Periodic (orange). The contours of the 2D probability densities are chosen to include 90% of the probability.
@@
Figure 2: Corner plot of the posteriors of the different models. DRW is blue, modified DRW is red, QPO is green, and Periodic is orange. Contours are drawn at a level to include 90% of the posterior probability. The black dashed lines refer to the period reported in Fatović et al. (2025). For reasons of clarity, although the fit is performed on standardized data, we report the marginal posterior distribution of the period in units of $\log_{10}$ Days.
@@
Combining the results shows in Fig. 1, Fig. 2, and Tab. 1, allows us to highlight some key insights. Visually, little difference is observed between a DRW (Fig. 1-a) and a Modified DRW (Fig. 1-b). The Bayes factor between the two models is close to 1, indicating no strong evidence in favour of any of the two models. We stress that the DRW model corresponds to the Modified DRW with $\gamma=1$. In principle, instead of considering $\gamma$ as a free parameter, we could have fixed it to a value different from 1 (e.g., $\gamma\simeq 1.5$). Such a model would have resulted in an only slightly larger evidence (i.e. $\log{Z}\simeq 138.5$), still not indicating a strong preference for either of the two models. Nevertheless, the comparison is useful: the estimated scale length and dispersion are shifted to smaller and larger values, respectively. This indicates that the estimates of $\tau$ and $\sigma$ parameters depend on the high-frequency slope of the assumed power spectrum, possibly suggesting that $\gamma$ might be included to provide better estimates of the physical parameters (e.g., black hole mass and accretion rate, Arévalo et al., 2023, 2024) regulating the light-curve.
@@
The QPO model exhibits an interesting behaviour. As shown in Fig. 2, its posterior distribution peaks at values of $\tau$ and $\sigma$ that are highly compatible with the ones obtained with the DRW model. However, the period $T$ not only shows a peaked distribution corresponding to $\log_{10}{T}\simeq-1.0$ (i.e., $T\simeq 290$ days), but features a very pronounced tail extending toward larger values. This result suggests that, although the peak at $\log_{10}{T}\simeq-1.0$ is well-defined and sharp, the majority portion of the cumulative distribution lies at longer periods. It follows that for values of $T$ comparable or longer than the observational time span, the QPO model becomes insensitive to the oscillatory terms. From a Bayesian statistics perspective, it is not a problem: in cases where a strong periodic signal is present in data, models with ineffective oscillatory terms would contribute negligibly to the posterior and would not influence the final estimate of the best-fit model. However, that is not the case for J2320+0024. Indeed, as shown in Fig. 1-c, the shape of the QPO posterior distribution is also responsible for the difference between the maximum likelihood and the median model.
@@
The red curve demonstrates fast oscillations (i.e., $\log_{10}{T}\simeq-1\simeq 290$ days) while the blue line is much closer to a DRW model. Moreover, the Bayes ratio does not reveal any significant preference between the (modified) DRW and the QPO models, preventing any strong claim in favour of the latter. These findings suggest a lack of robust evidence for the presence of a periodic signal in the light-curve of J2320-0024.
@@
To compare our results with those found in Fatović et al. (2025) and to assess the potential impact of the overfitting on the QPO model, we also provide a fit with a Periodic model. We note that, unlike the other models considered in this work, the evidence of the Periodic model is significantly smaller. This finding clearly suggests that, analysing the current data, there is no statistical support for preferring the Periodic model over the alternative ones. Thus, there is no strong evidence that the light-curve of J2320+0024 can be described by a purely periodic signal, underlying the importance of red-noise intrinsic variability. Interestingly, the marginalised posterior distribution reveals extremely sharp peaks in the period ($T\simeq 1100$ days) with relative errors smaller than a few percent. Notably, the only peak we find in the period of the Periodic model does not match the main peak identified by the QPO model; instead, it overlaps with a secondary and smaller peak. It is important to notice that the nested sampling algorithms used with the Periodic model detect other peaks, including one that corresponds to the periodicity reported in Fatović et al. (2025). However, these smaller peaks have significantly lower likelihoods and are therefore rejected by our approach. Since in Fatović et al. (2025) a maximum likelihood approach is adopted and the period search is initialised close to a period of $T\simeq 260$ days, it might be possible that the period the authors found in that work corresponds to a local and not a global maximum of the likelihood. Similar conclusions can be drawn by analysing the histogram of the residuals (see Appendix B) among the four models. Although the Kolmogorov-Smirnov test (Smirnov, 1948) indicates that the residuals of the DRW and Modified DRW model always resemble a Gaussian distribution (i.e., p-value $>0.05$), this is not true in the case of the Periodic model, and only partially true (i.e., only when using the best-fit model) for the QPO one.
4 Discussion & Conclusion
@@
J2320+0024 has been proposed as a candidate for hosting a MBHB based on modulation of its light-curve (Fatović et al., 2023). However, the results we obtain following a solid Bayesian approach do not find evidence of a periodic signal in the data. This highlights the importance of preferring statistically robust methods when assessing the nature of variability present in a given source. Nevertheless, the absence of a periodic signal detection in the light-curve of J2320+0024 does not necessarily imply that a MBHB is not present in this source. Our result suggests that, even if such a periodic signal exists, the current data set is too limited in coverage and in robustly detecting it.
@@
Still, we also found a peak in the period for the QPO and Periodic model, which, once translated into physical units, correspond to $T=290$ days and $T=1100$ days respectively. Fatović et al. (2025) reported a periodicity of $T\simeq 270$ days, much closer to that found by our QPO model. We note that a period of $1100$ days is almost an integer multiple of $T=270$ days (i.e., close to four times larger). Interestingly, when our analysis is extended to the r-band (i.e., the same reported in Fatović et al., 2025), the detection of the short timescale periodicity (i.e., $T=290$ days) disappears, while the identification of a longer, statistically unfavoured, peak (i.e., $T\simeq 1100$ days) persists. Such changes in the estimation of the period between adjacent broadbands further support the lack of a robust signal in the light-curve of J2320+0024.
@@
We believe that J2320+0024 is a very interesting source that deserves further follow-up observations and analysis, in particular to better characterise the spectral variability reported in Fatović et al. (2025). The authors explained this behaviour within the framework of the PoSKI model (Popović et al., 2021). This model assumes the presence of a MBHB surrounded by a common BLR, illuminated by the emission of the two mini-discs around each binary component. In this scenario, the appearance of a variable blue or red wing in the MgII is attributed to the binary orbital motion, and it is expected to mirror the periodicity observed in the light-curve. Although this scenario is certainly intriguing, our results challenge it. Indeed, alternative explanations for the observed MgII variability are possible. As discussed in the Sec. 1, several physical mechanisms, such as non-symmetric BLR (Rigamonti et al., 2025), can produce distorted broad emission profiles, possibly explaining the features observed in the MgII line of J2320+0024. However, explaining the observed variation over time is more challenging. We note that the most significant variation occurs between the SDSS spectrum, obtained about 8 years ago, and the more recent Magellan or Gemini spectra. Over such long timescales, comparable with the BLR dynamical timescales, variations in the shape of a BEL might be expected, possibly as a result of changes in the morphology or dynamics of the overall BLR. Moreover, the more recent Magellan and Gemini spectra show slight variations among each other; however, those are comparable with the variation of a non-symmetric BLR reverberating after a DRW-like variability of the AGN (Sottocorno et al., 2025).
@@
We still stress that further photometric and spectroscopic follow-up observations are needed to better clarify the nature of J2320+0024. In particular, as mentioned in Sec. 1, simultaneous modelling of repeated photometric and spectroscopic observations could unveil the true nature of this source. For instance, the light-curve modelling presented here could be extended with a consistent response from perturbed a BLR (e.g., Rigamonti et al., 2025; Sottocorno et al., 2025) to provide a detailed comparison with modelling or tests based on the MBHB scenario (Popović et al., 2021; Bertassi et al., 2025).
Acknowledgements
FR acknowledges the support from the Next Generation EU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 - From Research to Business (M4C2), Investment Line 3.1 - Strengthening and creation of Research Infrastructures, Project IR0000012 – “CTA+ - Cherenkov Telescope Array Plus.
@@ Acknowledgements
MD acknowledge funding from MIUR under the grant PRIN 2017-MB8AEZ, financial support from ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU, and support by the Italian Ministry for Research and University (MUR) under Grant ’Progetto Dipartimenti di Eccellenza 2023-2027’ (BiCoQ). We acknowledge a financial contribution from the Bando Ricerca Fondamentale INAF 2022 Large Grant, Dual and binary supermassive black holes in the multi-messenger era: from galaxy mergers to gravitational waves. RB acknowledges support from the ICSC National Research Center funded by NextGenerationEU, and the Italian Space Agency grant Phase A activity for LISA mission, Agreement n.2017-29-H.0. The data underlying this article will be shared on reasonable request to the corresponding author. AS acknowledges the financial support provided under the European Union’s H2020 ERC CoG ”B Massive” (Grant Agreement: 818691) and AdG ”PINGU” (Grant Agreement: 101142079)
@@ Appendix A Analysis on the r-band
In this section, we repeat the same analysis of Sec. 3 on the r-band data of J2320+0024. Also in this case, we tested for different models: DRW, Modifiead DRW, QPO, and Periodic reporting their best-fit light-curves and parameters in Fig. 3 and Tab. 2 respectively [^7] Also in this case, we standardized the data using a mean r-band magnitude of 21.182 with a standard deviation of 0.319 and a mean modified Julian date (MJD) of 57602.992 with a standard deviation of 2650.236..
The overall considerations from the analysis are similar to those discussed for the g-band. The Periodic model is strongly unfavoured, while the DRW, Modified DRW, and QPO have comparable evidence. Notably, in this case, the Bayes ratios between these three models point toward a slight preference for the Modified DRW. Interestingly, while the period found in the Periodic model remains close to that found in the g-band, the main peak in the period of the QPO model is larger and no longer consistent with the one reported by Fatović et al. (2025) and found by us in the g-band (see Figs. 24). Such a large variation in the measured periodicities among adjacent bands further probes the lack of a robust periodicity detection in J2320+0024
Table 2: Summary of the different model parameters for the fit to the r-band light-curve of SDSS J2320+0024.
@@
Scale length $\log_{10}{\tau}$ $[-3,1.7]$ $-0.8^{+0.3}_{-0.2}$ $-1.3^{+0.1}_{-0.1}$ $-0.8^{+0.4}_{-0.3}$
Dispersion $\log_{10}{\sigma}$ $[-3,1.7]$ $0.5^{+0.1}_{-0.1}$ $0.7^{+0.1}_{-0.1}$ $0.5^{+0.1}_{-0.1}$ $0.01^{+0.3}_{-0.2}$
Slope $\gamma$ $[0.5,5]$ $2.0^{+0.2}_{-0.4}$
Period $\log_{10}{T}$ $[-2,2]$ $0.5^{+1}_{-0.7}$ $-0.395^{+0.001}_{-0.001}$
@@
Scale length $\log_{10}{\tau}$ $[-3,1.7]$ $-0.90$ $-1.30$ $-0.73$
Dispersion $\log_{10}{\sigma}$ $[-3,1.7]$ $0.51$ $0.74$ $0.44$ $-0.06$
Slope $\gamma$ $[0.5,5]$ $2.1$
Period $\log_{10}{T}$ $[-2,2]$ $-0.20$ $-0.39$
@@
 [^8] From left to right, the columns provide a brief description of each parameter, its reference name as used in this work, the assumed prior range, and the best-fit values with their credibility intervals for all the different models. The first table block reports the parameters estimated as the median, 16th, and 84th percentiles of the posterior distribution, while the second table block reports the set of parameters that maximize the likelihood. All the priors are log-uniform, and all the parameters are in dimensionless units since the data have been standardized before fitting. The last row reports the Bayesian log evidence of each model clearly disfavoring the Periodic model.
@@
Figure 3: Best-fit for the r-band light-curve of J2320+0024. The black points represent the data with their errors, the blue solid line and the shaded area represent the median, 16th, and 84th percentiles from the posterior distributions, and the red dashed line refers to the maximum likelihood model. From top to bottom: DRW (a), modified DRW (b), QPO (c), and Periodic (d), we refer to Tab. 1 for the estimated best-fit parameters.

### v2

@@ 3.2 Fit and model selection
Since our goal was to estimate the Bayesian evidence of the proposed models, we adopted a nested sampling algorithm for parameter inference (Skilling, 2004). In particular, following previous works (e.g. Rigamonti et al., 2022, 2023), we employed the python-based implementation raynest (Veitch et al., 2017), which efficiently computes the marginal likelihood (Eq. 1) using the GPyTorch package (see Gardner et al., 2021) [^5] While Eq. (1) is a marginal-likelihood for a single GP, it is instead to be interpreted as a likelihood when exploring GPs with different parameters.. We performed a parameter estimation with 1000 live points, assuming log-uniform priors for all parameters, except $\gamma$, which we assumed to have a uniform prior. The prior ranges adopted for all models are reported in Tab. 1.
@@
Figure 1: Best-fit to the g-band light curve of J2320+0024. Black points represent the data with their errors, the solid blue line and shaded area show the median, 16th, and 84th percentiles of the posterior distributions, and the red dashed line indicates the maximum-likelihood model. From top to bottom: DRW (a), modified DRW (b), QPO (c), and Periodic (d) . See Table 1 for the estimated best-fit parameters.
@@
Fig. 1 shows the best-fit light curves with associated errors for the four models discussed in Sec. 3.1: DRW (a), modified DRW (b), QPO (c), and periodic (d). In all plots, we report the data with their associated uncertainties; the posterior predictive median, 16th, and 84th percentiles; and the maximum likelihood model. More precisely, the posterior predictive median and percentiles were computed by sampling 100 GP realisations for a fixed set of posterior-sampled kernel hyperparameters. This allowed effective marginalisation over the kernel hyperparameters, preserving and propagating all information in their posterior distribution. The best-fit model parameters (either percentiles or likelihood maxima), together with their corresponding Bayesian evidence values, are reported in Tab. 1. It is important to clarify that, while the maximum likelihood models (i.e. dashed red lines) are obtained using the ”maximum likelihood” parameters reported in Tab. 1, this is not the case for the median, 16th, and 84th percentile models. In agreement with our Bayesian approach, the blue lines and the shaded areas presented in Fig. 1 represent the median, 16th, and 84th percentiles computed by evaluating the GP model on all parameter combinations describing the posterior distribution. Therefore, the blue shaded area reported in Fig. 1 encapsulates and propagates three different uncertainties: data variance (i.e. measured errors), kernel variance (i.e. the kernel value at $t_{i}=t_{j}$), and the posterior distribution of the kernel parameters. The combination and propagation of these three sources of error in the posterior predictive are highly non-linear and difficult to disentangle. For instance, measurement errors contribute to the production of a smoother model, less dependent on outliers and highly uncertain points, while the kernel variance and the hyperparameter uncertainty simultaneously contribute to the specific shape of the posterior predictive. These effects depend on the chosen model. For example, the periodic model exhibits a larger uncertainty in the posterior predictive with respect to the others, due to a combination of factors. Specifically, this model, unlike the others, fits the data poorly and therefore requires a larger $\sigma$. The large $\sigma$ is needed due to the specific shape of the covariance matrix (see Eq. 5), which suppresses kernel variance in the regions where the cosine is close to zero.Fig. 2 displays a comparison of the marginalised posterior distribution of the DRW (blue), modified DRW (red), QPO (green), and periodic (orange). The contours of the 2D probability densities were chosen to include 90% of the probability.
@@
Figure 2: Corner plot of the posteriors for the different models. The DRW model is shown in blue, modified DRW in red, QPO in green, and Periodic in orange. Contours are drawn to include 90% of the posterior probability. The dashed black lines indicate the period reported in Fatović et al. (2025). For clarity, although the fit is performed on standardised data, the marginal posterior distribution of the period is shown in units of $\log_{10}$ days.
@@
Combining the results shown in Fig. 1, Fig. 2, and Tab. 1, allows us to highlight some key insights. Visually, little difference is observed between the DRW (Fig. 1-a) and the modified DRW (Fig. 1-b) models. The Bayes factor between the two models is close to 1, indicating that there is no strong evidence in favour of either model. We stress that the DRW model corresponds to the modified DRW model with $\gamma=1$. In principle, instead of considering $\gamma$ as a free parameter, we could have fixed it to a value different from 1 (e.g. $\gamma\simeq 1.5$). Such a model would produce only slightly larger evidence (i.e. $\log{Z}\simeq 138.5$) and would still not indicate a strong preference for either model. Nevertheless, the comparison is useful: the estimated scale length and dispersion are shifted to smaller and larger values, respectively. This indicates that estimates of the $\tau$ and $\sigma$ parameters depend on the high-frequency slope of the assumed power spectrum, possibly suggesting that $\gamma$ might be included to provide better estimates of the physical parameters (e.g. black hole mass and accretion rate, Arévalo et al., 2023, 2024) that regulate the light curve.
@@
The QPO model exhibits an interesting behaviour. As shown in Fig. 2, its posterior distribution peaks at values of $\tau$ and $\sigma$ highly compatible with those obtained from the DRW model. However, the period $T$ not only shows a peaked distribution corresponding to $\log_{10}{T}\simeq-1.0$ (i.e. $T\simeq 290$ days), but also features a pronounced tail extending to larger values. This suggests that although the peak at $\log_{10}{T}\simeq-1.0$ is well defined and sharp, the majority of the cumulative distribution lies at longer periods. Consequently, for values of $T$ comparable or longer than the observational timespan, the QPO model becomes insensitive to the oscillatory terms. From a Bayesian statistics perspective, this is not problematic: when a strong periodic signal is present, models with ineffective oscillatory terms contribute negligibly to the posterior and do not influence the final best-fit model estimate. However, this is not the case for J2320+0024. Indeed, as shown in Fig. 1-c, the shape of the QPO posterior distribution is also responsible for the difference between the maximum likelihood and median models.
@@
The red curve demonstrates fast oscillations (i.e. $\log_{10}{T}\simeq-1\simeq 290$ days), whereas the blue line is much closer to a DRW model. Moreover, the Bayes ratio reveals no significant preference between the (modified) DRW and the QPO models, precluding strong claims in favour of the latter. These findings suggest a lack of robust evidence for a periodic signal in the light curve of J2320-0024.
@@
To compare our results with those found in Fatović et al. (2025) and to assess the possible impact of overfitting on the QPO model, we also fit a periodic model. We note that, unlike the other models considered in this work, the evidence for the periodic model is significantly lower. This clearly suggests that there is no statistical support for preferring the periodic model over the alternatives given the current data. Thus, there is no strong evidence that the light curve of J2320+0024 can be described by a purely periodic signal, highlighting the importance of red-noise intrinsic variability. Interestingly, the marginalised posterior distribution reveals extremely sharp period peaks ($T\simeq 1100$ days) with relative errors smaller than a few percent. Notably, the only peak in the period of the periodic model does not match the main peak identified by the QPO model; rather, it overlaps with a secondary, smaller peak. Nested sampling algorithms used with the periodic model detect other peaks, including one that corresponds to the periodicity reported in Fatović et al. (2025). However, these smaller peaks have significantly lower likelihoods and are rejected by our approach. Since Fatović et al. (2025) adopt a maximum likelihood approach with the period search initialised near $T\simeq 273$ days, the reported period may correspond to a local rather than a global likelihood maximum. Similar conclusions arise from analysing the histogram of the residuals (see Appendix C) across the four models. Although the Kolmogorov-Smirnov test (Smirnov, 1948) indicates that the residuals for the DRW and modified DRW models always resemble a Gaussian distribution (i.e., p-value >0.05>0.05), this is not the case for the periodic model and is only partially true for the QPO model (i.e. only when using the best-fit model).
4 Discussion and conclusion
@@
J2320+0024 has been proposed as a candidate hosting an SBHB based on modulation of its light curve (Fatović et al., 2023), who report a period of 278 days based on a Lomb-Scargle periodogram analysis of the g, r, and i bands, exceeding the false alarm probability threshold. However, our results, obtained via a rigorous Bayesian approach, find no evidence for a periodic signal in the data. This highlights the importance of applying statistically robust methods to assess the nature of variability in a given source and the exposes the limitations of periodogram analyses for detecting true periodicity in the presence of red noise. Nevertheless, the absence of a periodic signal in the light curve of J2320+0024 does not necessarily rule out the presence of an SBHB in this source. Our result suggests that even if such a periodic signal exists, the current dataset lacks sufficient coverage to detect it robustly.
@@
We also find period peaks for the QPO and periodic models that correspond to $T=290$ days and $T=1100$ days, respectively, when translated into physical units. Fatović et al. (2023, 2025) report a periodicity of $T\simeq 270$ days, much closer to the value of our QPO model. We note that a period of $1100$ days is almost an integer multiple of $T=270$ days (i.e. close to four times larger). Interestingly, when our analysis is extended to the r-band (i.e. the same reported in Fatović et al., 2025), the short timescale periodicity ($T=290$ days) disappears, while the longer, statistically unfavoured, peak at $T\simeq 1100$ days persists. Different behaviours are also observed in the i-band, where the short timescale periodicity of the QPO is not detected, and a double peak appears in the periodic model. These variations in period estimation between adjacent broad bands further support the absence of a robust signal in the light curve of J2320+0024.
@@
We consider J2320+0024 a very interesting source deserving further follow-up observations and analysis, particularly to better characterise the spectral variability reported in Fatović et al. (2025). This behaviour was explained within the framework of the Popović, Simić, Kovačević, Ilić (PoSKI) model model (Popović et al., 2021). The model assumes the presence of an SBHB surrounded by a common circumbinary BLR, illuminated by emission from the two mini-discs around each binary component, with only the less massive black hole retaining its own BLR. In this scenario, the variable blue or red wing observed in MgII is attributed to binary orbital motion, particularly to Doppler-boosted emission from the less massive black hole’s BLR, and is expected to reflect the periodicity observed in the light curve. While intriguing, our results challenge this scenario. Indeed, as Fatović et al. (2025) note, alternative explanations for the observed MgII variability exist. As discussed in Sec. 1, physical mechanisms, such as non-symmetric BLR (Storchi-Bergmann et al., 2003; Pancoast et al., 2014; Raimundo et al., 2020; Rigamonti et al., 2025), can produce distorted broad emission profiles, possibly explaining the features observed in the MgII line of J2320+0024. However, explaining the observed variation over time remains more challenging. The most significant variation occurs between the SDSS spectrum, obtained approximately years ago, and the more recent Magellan or Gemini spectra. Over such long timescales, comparable to BLR dynamical timescales, variations in BEL shape might be expected, possibly due to changes in the morphology or dynamics of the overall BLR. Furthermore, the more recent Magellan and Gemini spectra show exhibit variations among themselves, comparable to variations from a non-symmetric BLR reverberating after DRW-like variability of the AGN (Sottocorno et al., 2025).
@@
We stress that further photometric and spectroscopic follow-up observations are necessary to clarify the nature of J2320+0024. In particular, as mentioned in Sec. 1, simultaneous modelling of repeated photometric and spectroscopic observations could reveal the true nature of this source. For example, the light curve modelling presented here could be extended to include a consistent response from a perturbed BLR (e.g. Rigamonti et al., 2025; Sottocorno et al., 2025) to provide a detailed comparison with models or tests based on the SBHB scenario (Popović et al., 2021; Bertassi et al., 2025).
Acknowledgements.
@@
We thank the anonymous referee for their comments and suggestions that helped us to improve the quality of the paper. FR acknowledges the support from the Next Generation EU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 - From Research to Business (M4C2), Investment Line 3.1 - Strengthening and creation of Research Infrastructures, Project IR0000012 – “CTA+ - Cherenkov Telescope Array Plus. MD acknowledge funding from MIUR under the grant PRIN 2017-MB8AEZ, financial support from ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU, and support by the Italian Ministry for Research and University (MUR) under Grant ’Progetto Dipartimenti di Eccellenza 2023-2027’ (BiCoQ). We acknowledge a financial contribution from the Bando Ricerca Fondamentale INAF 2022 Large Grant, Dual and binary supermassive black holes in the multi-messenger era: from galaxy mergers to gravitational waves. RB acknowledges support from the ICSC National Research Center funded by NextGenerationEU, and the Italian Space Agency grant Phase A activity for LISA mission, Agreement n.2017-29-H.0. The data underlying this article will be shared on reasonable request to the corresponding author. AS acknowledges the financial support provided under the European Union’s H2020 ERC CoG ”B Massive” (Grant Agreement: 818691) and AdG ”PINGU” (Grant Agreement: 101142079)
@@ Appendix A Analysis on the r-band
In this section, we repeat the same analysis of Sec. 3 on the r-band data of J2320+0024. Also in this case, we tested four different models: DRW, Modified DRW, QPO, and Periodic reporting their best-fit light curves and parameters in Fig. 3 and Tab. 2 respectively [^6] Also in this case, we standardised the data using a mean r-band magnitude of 21.182 with a standard deviation of 0.319 and a mean modified Julian date (MJD) of 57602.992 with a standard deviation of 2650.236..
The overall considerations from the analysis are similar to those discussed for the g-band. The Periodic model is strongly unfavoured, while the DRW, Modified DRW, and QPO have comparable evidence. Notably, in this case, the Bayes ratios between these three models point toward a slight preference for the Modified DRW. Interestingly, while the period found in the Periodic model remains close to that found in the g-band, the main peak in the period of the QPO model is larger and no longer consistent with the one reported by Fatović et al. (2025) and found by us in the g-band (see Figs. 2, 4). Such a large variation in the measured periodicities among adjacent bands further probes the lack of a robust periodicity detection in J2320+0024
Table 2: Summary of the different model parameters for the fit to the r-band light curve of SDSS J2320+0024.  [^7] From left to right, the columns provide a brief description of each parameter, its reference name as used in this work, the assumed prior range, and the best-fit values with their credibility intervals for all the different models. The first table block reports the parameters estimated as the median, 16th, and 84th percentiles of the posterior distribution, while the second table block reports the set of parameters that maximise the likelihood. All the priors are log-uniform, and all the parameters are in dimensionless units since the data have been standardised before fitting. The last row reports the Bayesian log evidence of each model clearly disfavoring the Periodic model.
@@
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] $0.8^{+0.3}_{-0.2}$ $1.3^{+0.1}_{-0.1}$ $0.8^{+0.4}_{-0.3} -$
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.5^{+0.1}_{-0.1}$ $0.7^{+0.1}_{-0.1}$ $5^{+0.1}_{-0.1} 0.0$$01^{+0.3}_{-0.2}$
Slope $\gamma$ $[0.5,5]$ - $2.0^{+0.2}_{-0.4}$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - $0.5^{+1}_{-0.7}$ $0.395^{+0.001}_{-0.001}$
@@
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] 0.90-0.90 1.30-1.30 0.73-0.73 -
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.51$ $0.74$ $0.44$ 0.06-0.06
Slope $\gamma$ $[0.5,5]$ - $2.1$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - 0.20-0.20 0.39-0.39
@@
Figure 3: Best-fit for the r-band light curve of J2320+0024. The black points represent the data with their errors, the blue solid line and the shaded area represent the median, 16th, and 84th percentiles from the posterior distributions, and the red dashed line refers to the maximum likelihood model. From top to bottom: DRW (a), modified DRW (b), QPO (c), and Periodic (d), we refer to Tab. 1 for the estimated best-fit parameters.